OpenStudy (anonymous):

Am I right? Multiply. (4x + 7)2 16x2 + 49 (My Answer) 16x2 + 28x + 49 16x2 + 56x + 49 8x2 + 14

OpenStudy (austinl):

That isn't what I get for an answer. Recall, \((ax+b)c=(a\times c)x+bc\)

OpenStudy (shamil98):

No, you forgot that (4x+7)(4x+7) 4x x 7 and 7 x 4x.

OpenStudy (shamil98):

which is 28x + 28x..

OpenStudy (austinl):

Or is it, \((4x+7)^2\) ?

OpenStudy (anonymous):

Yes that's what it is

OpenStudy (shamil98):

The answer is c... you would only get your answer if the values were (4x-7)^2

OpenStudy (anonymous):

That is the value

OpenStudy (anonymous):

are you multiplying it by two or are you given the power to two????

OpenStudy (austinl):

Ok, \((ax+b)^2=(ax+b)(ax+b)\Rightarrow~ax^2+abx+abx+b^2\)

OpenStudy (anonymous):

power of 2

OpenStudy (anonymous):

And so are the answer chocies

OpenStudy (anonymous):

@austinL is right

OpenStudy (anonymous):

so is it 16x2 + 28x + 49

OpenStudy (shamil98):

No.

OpenStudy (anonymous):

So then 16x2 + 56x + 49

OpenStudy (austinl):

So from your question of, \((4x+7)^2=(4x+7)(4x+7)\Rightarrow~(4x\times4x)+(7\cdot4)x+(7\cdot4)x+(7\cdot7)\)

OpenStudy (austinl):

When simplified, \(16x^2+56x+49\) That is correctomundo!

OpenStudy (anonymous):

Okay! Thanks for explaining! ^-^

OpenStudy (austinl):

No problem!